Sue and Neal are twins. Sue invests $5,000 at 7 percent when she is 25 years old. Neal invests $7,000 at 7 percent when he is 30 years old. Both investments compound interest annually. Both Sue and Neal retire at age 60. Who will have more money in their accounts at the time of retirement?
Answer #1

Sue money in the account at the time of retirement=PV*(1+r)^n

=5000*(1+7%)^35

=53382.91

Neal money in the account at the time of retirement=PV*(1+r)^n

=7000*(1+7%)^30

=53285.79

so from above Sue will have more money at the time of retirement

the above is answer..